# A Generalized Theory for Optical Cooling of a Trapped Atom with Spin

### v2

@@ II.2 Effective master equation within ground-state subspace
The effective spin operators are valid for any ground and excited spin states. The excited state only affects the $C^{(i)}$ coefficients.
The ground-state collapse operators in Eqs. 18 (far detuned limit) and 14 (near-resonant limit) describe the redistribution of the population within the ground-state. With the effective ground-state Hamiltonian, a master equation incorporating these operators describes the ground-state dynamics. The Hamiltonian accounts for energy shifts and coherent transfer between states, while the collapse operators account for incoherent redistribution due to spontaneous emission.
Next, we will derive an approximate population rate equation by working in the basis where $H_{\text{eff}}$ is diagonal. We define an interaction picture by applying a unitary transformation $\tilde{A}=UAU^{\dagger}$ with $U=e^{-iH_{\text{eff}}t}$. In the interaction picture, the time evolution of the Hamiltonian is included within the density matrix and operators $U\rho U^{\dagger}$ with unitary $U=e^{-iHt}$. In the interaction picture, the master equation is:
@@
Finally, we derive rate equations for the populations by multiplying by $\langle n|$ on the left and $|n\rangle$ on the right and defining the populations $P(n)=\tilde{\rho}_{nn}=\rho_{nn}$. The resulting equation contains coherence terms $\langle n|\rho|m\rangle$ between ground-states. If we work in the basis where the ground-state Hamiltonian is diagonal, these coherences are less important and can often be ignored. By setting all these coherences to zero, we obtain a population rate equation:
@@
where the transfer rate between ground-states $n$ and $m$ is:
@@
Note that in this last equation, $L_{\text{eff}}$ can be taken out of the interaction picture because as long as $n$ and $m$ are eigenstates of $H_{\text{eff}}$, the time-dependence cancels upon taking the magnitude. However, it is important to remember that the eigenstates must be those of $H_{\text{eff}}$. The steady-state conditions can then be found by solving the matrix equation for $\frac{dP(n)}{dt}=0$. In this paper, we derive analytic solutions for different cooling schemes by picking an appropriate basis where the coherences between eigenstates are least important and solving this rate equation.
@@ III Sideband cooling
Figure 2: (a) Resolved ($\Gamma<\nu$) and unresolved ($\Gamma>\nu$) sideband cooling in the Lamb-Dicke regime. (b) The cooling laser drives the three blue transitions. Spontaneous emission (red) most likely does not change $n$ in the LD regime. (c-d) Simulation of $\langle n\rangle$ in the unresolved and resolved regime for $\eta=0.05$. (e) The population versus trapping frequency. (f) Cooling in a standing wave can lower the energy by positioning the atom at the node, thereby decreasing scattering.
For an atom confined in a harmonic trap within the Lamb-Dicke (LD) regime, Doppler cooling operates differently than in free space, as illustrated in Fig. 2(a). In this scenario, the cooling laser drives transitions from the motional state $n$ to $n-1$, while spontaneous emission predominantly returns the atom to the $n-1$ state. This process leads to net cooling because, in the LD regime, the recoil energy from photon emission is smaller than the energy spacing between trap levels ($\hbar\nu$). Doppler cooling of trapped atoms can be classified into two regimes based on the relationship between the trapping frequency $\nu$ and the atomic transition linewidth $\Gamma$:
@@
These regimes exhibit distinct cooling dynamics and efficiencies, which we will explore in subsequent sections. Sideband cooling was extensively studied theoretically in the 1980’s by Refs. [36, 40, 30, 31]. It has been demonstrated experimentally in ions [29], neutral atoms [41, 42], and nanomechanical oscillators [28]. Here we present these results in our formalism. We then look at the case of different trapping frequencies and derive an analytic expression for the steady-state temperature.
Doppler cooling for a bound atom in the LD regime only requires a single traveling wave $\bm{\epsilon}(\hat{x})=e^{-ik\hat{x}}$. For a two-level system, the dipole operators are $\hat{D}=\hat{\sigma}=|g\rangle\langle e|$. The Hamiltonian is
$H=\nu\hat{a}^{\dagger}\hat{a}+\Delta|e\rangle\langle e|-\frac{\Omega}{2}\left(e^{-ik\hat{x}}\hat{D}^{\dagger}+h.c.\right),$ (25)
and the collapse operators for emission into the positive and negative $x$-direction is
@@
The resonant and spontaneous emission processes are shown in Fig. 2(b). To estimate the final population distribution, we can use the effective ground-state collapse operator in Eq. 14 and the population transfer rates from Eq. 24. From these equations, the transfer rate between ground-states $n_{1}$ and $n_{2}$ in the ground-state through the excited harmonic states $n_{e}$ is
@@
The sum inside the magnitude is over the excited state $n_{e}$ and pathways through each excited motional state can interfere. The sum on the outside is over the emission direction. Emission into each direction results in a different phase, and the average over all directions averages out the coherences between the different pathways. As a result, we can take the sum over the excited states outside of the sum and omit the average over directions,
@@
This rate describes two processes. The first term in the numerator $|\langle n_{2}|e^{ik\hat{x}}|n_{e}\rangle|^{2}$ describes the process of recoil due to spontaneous emission. The rest of the expression gives the rate of absorption and recoil from photons from the cooling light. Together, these two terms give the total rate of photon absorption and emission into motional states.
@@
$\gamma_{n\rightarrow n-1}=\eta^{2}(n)\,A_{\scriptscriptstyle-}$ (30)
@@
$A_{\scriptscriptstyle-}=R(\Delta+\nu)+\alpha R(\Delta).$ (32)
@@
Because the population only transfers between neighboring $n$ in the LD regime, we can calculate the steady-state population by assuming no flow between two states: $P(n+1)\,\gamma_{n+1\rightarrow n}=P(n)\,\gamma_{n\rightarrow n+1}$. Solving this equation leads to the steady-state condition
@@
$\frac{P(n+1)}{P(n)}=\exp{(-\hbar\nu/k_{B}T)}.$
@@
Multiplying by $n$ and summing over $n$ gives the time rate equation for $\langle n\rangle$:
@@
The cooling rate is $W_{A}=\eta^{2}(A_{\scriptscriptstyle-}-A_{\scriptscriptstyle+})$, which is balanced by the constant heating from spontaneous emission $\eta^{2}A_{\scriptscriptstyle+}$. The steady-state solution is
@@
In the resolved-sideband regime ($\Gamma\ll\nu$), the lowest temperature occurs at the cooling sideband $\Delta=-\nu$, at which the steady-state is
@@
In the unresolved-sideband regime ($\Gamma\gg\nu$), the steady state is [36, 30]
@@
In Fig. 2 we simulate the exact solution using the master equation from Eq. 3. Fig. 2(c-d) are in the unresolved and resolved sideband regimes for $\eta=0.05$ and in the low power limit $\Omega=0.05/\Gamma$. In the resolved regime in (d), best cooling occurs at $\Delta=-\nu$ and agrees closely with Eq. 38. The unresolved regime in (e) is a zoomed plot of the lower left corner of (d). Theoretically, the lowest temperature occurs for $\Delta=-\Gamma/2$, although the exact solution shows that it is broad and not much better than the resolved criteria $\Delta=-\Gamma$.
In plot Fig. 2(e), we plot steady-state population versus the trapping frequency for various LD parameters $\eta$. The population converges for $\eta<0.1$.
The cooling performance of sideband cooling can be improved by adding a second counter-propagating beam with the same polarization to form a standing wave and placing the atom at the node of the field [43]. Although the atom now sees zero intensity at the equilibrium position, the sidebands are still driven. This can be seen by expanding the atom-field interaction as a series to first order in terms of the LD parameter and field gradient
@@
The gradient can be either a phase or amplitude gradient, both of which drive the sideband. Intensity at the position of the atom does not lead to cooling and, in fact, only leads to extra heating. If we place the atom in a standing wave with linear polarization, the field is $\bm{\epsilon}(\hat{x})=\sin(k\hat{x})$. Fig. 2(f) simulates the steady-state population versus atom position with $\nu$ = $-\Delta$. Theoretically, this gets rid of the process in Fig. 2(b) of driving $n$ to $n$, followed by spontaneous emission heating to the sideband. The steady-state solution is the same except with $\alpha=0$. Cooling can be further enhanced by placing the atom in a cavity that can enhance the red sideband [44, 45].
III.1 Excited state with different trapping frequency
The Doppler cooling section assumed that the electronic ground and excited states have the same trapping frequency. But typically, the ground and excited states of neutral atoms have different polarizabilities except at “magic wavelengths.” The different potentials lead to heating due to the abruptly changing dipole force during absorption or emission, shown in the schematics in Fig. 3. In the extreme case, the excited state is anti-trapped resulting in loss of the atom if the excited lifetime is long compared to $1/\nu$.
Because magic wavelengths are not often convenient for optical trapping, neutral atoms more often use two-photon cooling schemes, such as polarization gradient cooling and Raman sideband cooling. However, alkaline-earth atoms contain narrow linewidth transitions that have been used for imaging and cooling [46, 10], even ground-state cooling in Ref. [47]. Alkali atoms also contain narrow lines from forbidden transitions like the quadrupole transition of the S to D transition [48] or (n)S to (n+1)P transitions [49].
@@ III.1 Excited state with different trapping frequency
We now derive the equations for an atom with a ground-state frequency $\nu$ and an excited state frequency $\nu_{e}$. The excited state harmonic eigenstates are stretched or compressed relative to the ground-state eigenstates. The squeezing operator from quantum optics can therefore be utilized to relate the wavefunctions in the two potentials:  [51, 52, 53]
@@
where $|n_{e}\rangle$ is the ground-state basis $|m_{e}\rangle$ is the excited state basis. The squeezing operator $\hat{S}(r)$ is
@@
The excited state annihilation operator is related to the ground-state annihilation operator by
@@
The harmonic Hamiltonian is
$H_{\text{ h}}$ $=\nu(\hat{a}^{\dagger}\hat{a})P_{g}+\nu_{e}(\hat{a}_{e}^{\dagger}\hat{a}_{e})P_{e}$
@@
Figure 3: Sideband cooling with different trapping frequencies. (a) Two regimes, where the excited state is less trapped (left) or more trapped (right) result in extra heating. This heating comes from coupling many different motional states in the quantum picture, or a fluctuating dipole force in the classical picture. For different trapping frequencies, the $n$ to $n-1$ transition can be resonant for only one specific $n$. For $\nu_{e}<\nu$, higher $n$ gradually shift more blue towards heating. For $\nu_{e}>\nu$, higher $n$ gradually shifts more red, leading to cooling of higher order sidebands. (b) Population for a different excited state trapping frequency $\nu_{e}$, plotted for different $\nu/\Gamma$ and $\eta=0.1$ and $\Delta=-\nu$. The magic condition is $\nu_{e}/\nu=1$. (c) Population versus detuning for different excited to ground trapping frequency ratios $\nu_{e}/\nu$. A minimum occurs at $\Delta=-\nu$, but a second minimum occurs at $\Delta=-2\nu$ when $\nu_{e}\neq\nu$. Plot is for $\eta=0.1$, $\nu/\Gamma=2$. Eq. 59 is plotted in a light dashed line. Different trapping frequencies allow the coupling of higher order sidebands, providing new opportunities for cooling.
The transfer rates between the ground-states are similar to Eq. 27, where now the squeezing operators are used to modify the excited harmonic states:
@@
The sum over the excited states cannot generally be taken out of the magnitude now because each term does not necessarily have the directional emission terms. The transition matrix elements now contain the squeezing operator. To first order in $r$ and $\eta$, the matrix elements are
@@
The first order LD terms couple $n\rightarrow n\pm 1$. In contrast, the first-order squeezing terms couple $n\rightarrow n\pm 2$, preserving the symmetry of the wavefunction.
The resonance condition for the light absorption is now determined by the detuning $\Delta+\nu n_{1}-\nu_{e}n_{e}$, which we can re-express in terms of the difference trapping frequency can be rewritten as $\Delta+\nu n_{1}-\nu_{e}n_{e}=(\Delta+n_{e}\Delta\nu)+\nu(\nu_{1}-\nu_{e})$, where the trapping frequency difference is $\Delta\nu=\nu_{e}-\nu$. For increasing $n$, the detuning between $n_{1}=n_{e}$ changes by $\nu n$. We define the resonance condition for the $n$’th harmonic state as
@@
$\gamma_{n\rightarrow n-1}=\eta^{2}\,n\,A_{\scriptscriptstyle-}.$ (50)
@@
$A_{\scriptscriptstyle-}=R({\Delta}_{n}+\nu)+\alpha R({\Delta}_{n}).$ (52)
In this context, the rates add incoherently due to the randomized phase acquired during emission in multiple directions. As illustrated in Fig. 3(a), the $n$-dependent detuning results in the first-order cooling sideband being resonant only for a specific vibrational quantum number $n$. For a less tightly trapped excited state (left panel), the cooling light shifts towards heating sidebands for higher $n$ values. Conversely, lower $n$ values gradually shift towards cooling of higher-order sidebands. This effect leads to instability when $\nu_{e}/\nu<1$, where higher-energy populations heat out of the trap. When $\nu_{e}/\nu>1$, a type of “cap” emerges, with heating occurring below a certain $n$ value and cooling above it. Consequently, this cap effectively traps population below it.
@@
$\gamma_{n\rightarrow n-2}=r^{2}\,n(n-1)\,\,B_{\scriptscriptstyle-}.$ (54)
The new rates $B$ are a sum over different excited states, but the coherence/interference now matters because the coupling does not involve spontaneous emission to first order,
@@
$B_{\scriptscriptstyle+}=\left|{T}({\Delta}_{n})-{T}({\Delta}_{n}-2\nu)\right|^{2}.$ (56)
@@
Cooling is a balance between the sideband and squeezing terms. Cooling due to sidebands is $W_{A}\approx\eta^{2}(A_{-}-A_{+})$, while squeezing cooling is $W_{B}\approx 2r^{2}(B_{-}-B_{+})$, with the squeezing dependence being quadratic. Additionally, there is heating from both that goes as $\eta^{2}A_{+}$ and $2r^{2}B_{+}$, respectively. Interestingly, the squeezing results in a coupling between different motional states and can be used for cooling, as shown below.
Fig. 3(b) is a master equation simulation of steady-state population versus the excited to ground-state trapping frequency ratio $\nu_{e}/\nu$. The magic condition is $\nu_{e}/\nu=1$. Cooling performance drops on both sides, with much more dramatic for a weaker excited state ($\nu_{e}/\nu<1$). Fig. 3(c) is a detuning scan for $\eta=0.1$, $\nu=2\Gamma$, and various $\nu_{e}/\nu$. There is a resonance at $\Delta=-\nu$ due to sideband cooling, but also a second resonance at $\Delta=-2\nu$ due to the squeezing coupling.
For cooling the atom in a ‘non-magic’ trap, a technique known as “Sisyphus Cooling”, has been recently demonstrated in tweezers [54, 55]. In situations where the polarizability of the excited state is greater (or less) than that of the ground-state, the atom experiences a deeper (or shallower) trap upon excitation by the cooling laser. This difference enables the implementation of ‘attractive’ (or ‘repulsive’) Sisyphus cooling. In the repulsive Sisyphus cooling method, there is a limit known as the Sisyphus cap to how much the atom can be cooled. However, in the attractive Sisyphus regime, significantly lower temperatures can be achieved within the trap. Ref. [50] proposes that sweeping the cooling laser’s frequency adiabatically could achieve temperatures well below the Sisyphus cap.

### v3

@@ II.2 Effective master equation within ground state subspace
The effective spin operators are valid for any ground and excited spin states. Interestingly, the excited state spin only affects the $C^{(i)}$ coefficients.
The ground state collapse operators in Eqs. 19 (far-detuned limit) and 15 (near-resonant limit) describe the redistribution of the population within the ground state. With the effective ground state Hamiltonian, a master equation incorporating these operators describes the ground state dynamics. The Hamiltonian accounts for energy shifts and coherent transfer between states, while the collapse operators account for incoherent redistribution due to spontaneous emission.
Next, we will derive an approximate population rate equation for the ground state spin levels by working in the basis where $H_{\text{eff}}$ is diagonal. While $H_{\text{eff}}$ captures the coherent dynamics within the ground state, the collapse operators capture population redistribution from spontaneous emission. We define an interaction picture by applying a unitary transformation $\tilde{A}=UAU^{\dagger}$ with $U=e^{-iH_{\text{eff}}t}$. In the interaction picture, the time evolution of the Hamiltonian is included within the density matrix and operators $U\rho U^{\dagger}$. In the interaction picture, the master equation is:
@@
Finally, we derive rate equations for the populations by multiplying by $\langle n|$ on the left and $|n\rangle$ on the right and defining the populations $P(n)=\tilde{\rho}_{nn}=\rho_{nn}$. The resulting equation contains coherence terms $\langle n|\rho|m\rangle$ between ground states. If we work in the basis where the ground state Hamiltonian is diagonal, these coherences are less important and can often be ignored. By setting all these coherences to zero, we obtain a population rate equation:
@@
where the transfer rate between ground states $n$ and $m$ is:
@@
Note that in this last equation, $L_{\text{eff}}$ can be taken out of the interaction picture because as long as $n$ and $m$ are eigenstates of $H_{\text{eff}}$, the time-dependence cancels upon taking the magnitude. However, it is important to remember that the eigenstates must be those of $H_{\text{eff}}$. The steady-state conditions can then be found by solving the matrix equation for $\frac{dP(n)}{dt}=0$. In this paper, we derive analytic solutions for different cooling schemes by first picking an appropriate basis where the coherences between eigenstates are least important and then solving this rate equation.
@@ III Sideband cooling
Figure 2: (a) Resolved ($\Gamma<\nu$) and unresolved ($\Gamma>\nu$) sideband cooling in the Lamb-Dicke regime. A red-detuned laser drives $n$ to $n-1$, which mostly decays down to the $n-1$ ground state. (b) The cooling laser drives the three blue transitions: $n\rightarrow n$ and $n\rightarrow n\pm 1$. Spontaneous emission (red) most likely does not change $n$ in the LD regime. (c-d) Simulation of $\langle n\rangle$ in the unresolved and resolved regime for a LD parameter $\eta=0.05$. (e) The final population versus trapping frequency for varying $\eta$. The lines converge below $\eta=0.1$. (f) Cooling in a standing wave can lower the energy by positioning the atom at the node of the field, thereby decreasing scattering but still allowing sideband cooling due to the presence of a gradient field.
In this section, we start looking at specific cooling schemes, starting with sideband cooling, a single-photon cooling schemes that works even for a two-level system. In free-space Doppler cooling, an atom moves towards a red-detuned light source, and the Doppler effect brings them closer to resonance, causing them to absorb photons and receive momentum kicks that oppose the motion.
For an atom confined in a harmonic trap within or close to the Lamb-Dicke (LD) regime, a different picture arises, as illustrated in Fig. 2(a). In this scenario, a red-detuned cooling laser drives transitions from the motional state $n$ to $n-1$, while spontaneous emission predominantly does not change $n$, returning the atom to the ground $n-1$ state. This process leads to net cooling because, in the LD regime, the recoil energy from photon emission is smaller than the energy spacing between trap levels ($\hbar\nu$).
Doppler cooling of trapped atoms can be classified into two regimes based on the relationship between the trapping frequency $\nu$ and the atomic transition linewidth $\Gamma$:
@@
These regimes exhibit distinct cooling dynamics and efficiencies. Sideband cooling was extensively studied theoretically in the 1980’s by Refs. [36, 42, 30, 31]. It has been demonstrated experimentally in ions [29], neutral atoms [43, 44], and nanomechanical oscillators [28]. Here we present these results in our formalism. Unlike for ion traps, optical traps typically have different potentials for the ground and excited states. We extend this formalism to treat different trapping frequencies and derive an analytic expression for the steady-state temperature as a function of the mismatch.
III.1 Sideband cooling for the same ground and excited trapping frequencies
Doppler cooling for a bound atom in the LD regime only requires a single traveling wave $\bm{\epsilon}(\hat{x})=e^{ik\hat{x}}$. For a two-level system, the dipole operators are $\hat{D}=\hat{\sigma}=|g\rangle\langle e|$. The Hamiltonian is
$H=\nu\hat{a}^{\dagger}\hat{a}+\Delta|e\rangle\langle e|-\frac{\Omega}{2}\left(e^{ik\hat{x}}\hat{D}^{\dagger}+h.c.\right),$ (26)
and the collapse operators for emission into the positive and negative $x$-directions are
@@ III.1 Sideband cooling for the same ground and excited trapping frequencies
The resonant and spontaneous emission processes are shown in Fig. 2(b). In LD regime, the population in the ground $n$ state is driven to excited $n$ and $n\pm 1$. The matrix element to $n-1$ is reduced by $\eta\sqrt{n}$ relative to same $n$. However, as shown in (b), the transition can be preferred if made resonant by detuning the laser by $\Delta=\nu$. The excited then decays back to the same $n$ with rate $\Gamma$, and changing $n$ with reduced rate $\Gamma\eta^{2}n$. To estimate the final population distribution, we can use the effective ground state collapse operator in Eq. 15 and the population transfer rates from Eq. 25. From these equations, the transfer rate between ground states $n_{1}$ and $n_{2}$ in the ground state through the excited states $n_{e}$ is
@@
The sum inside the magnitude is over all excited states $n_{e}$. Pathways from $n_{1}$ to $n_{2}$ ground states can in principle interfere through these excited states. The sum on the outside is over the emission direction. The random emission into different directions results in a random phase that ends up averaging out the coherence terms, suppressing the interference. As a result, we can take the sum over the excited states outside of the sum and omit the average over directions,
@@
This rate describes two processes. The first term in the numerator $|\langle n_{2}|e^{ik\hat{x}}|n_{e}\rangle|^{2}$ describes the process of recoil due to spontaneous emission. The rest of the expression gives the scattering rate of photon absorption. Together, these two terms give the total rate of photon absorption and emission into motional states and the resulting recoil.
@@
$\gamma_{n\rightarrow n-1}=\eta^{2}n\,A_{\scriptscriptstyle-}$
@@
$A_{\scriptscriptstyle-}=R(\Delta+\nu)+\alpha R(\Delta).$
@@
Next we calculate the steady-state temperature. Because the population only transfers between neighboring $n$ in the LD regime, we can calculate the steady-state population by assuming no flow of population between $n$ and $n+1$, which gives
$P(n+1)\,\gamma_{n+1\rightarrow n}=P(n)\,\gamma_{n\rightarrow n+1}.$
Solving this equation and substituting Eq. 30 leads to the steady-state condition
@@
$\frac{P(n+1)}{P(n)}=\exp{\left(-\frac{\hbar\nu}{k_{B}T}\right)}.$ (35)
@@
We can also derive an expression for the time evolution of the average population. Multiplying Eq. III.1 by $n$ and summing over $n$ gives the time rate equation for $\langle n\rangle$:
@@
The cooling rate $\eta^{2}(A_{\scriptscriptstyle-}-A_{\scriptscriptstyle+})$ is balanced by the constant heating from spontaneous emission $\eta^{2}A_{\scriptscriptstyle+}$. The steady-state solution is
@@
In the resolved-sideband regime ($\Gamma\ll\nu$), the lowest temperature occurs at the cooling sideband $\Delta=-\nu$, at which the steady-state is [36, 30]
@@
In the unresolved-sideband regime ($\Gamma\gg\nu$), the steady state is
@@
In Fig. 2, we simulate sideband cooling with the master equation from Eqs. 26 and 27. We use the steady-state solver in QuTip using up to 15 harmonic states in both the ground and excite state and assume 1D spontaneous emission. The simulations in Fig. 2(c-d) are in the unresolved and resolved sideband regimes for $\eta=0.05$ and in the low saturation limit $\Omega=0.05\,\Gamma$. In the resolved regime in (d), best cooling occurs at $\Delta=-\nu$ and agrees closely with Eq. 38. The unresolved regime in (c) is a zoomed in plot of the lower left corner of (d). Theoretically in the unresolved regime, the lowest temperature occurs for $\Delta=-\Gamma/2$, although the exact solution shows that the optimum is broad and not much better than the resolved criteria $\Delta=-\Gamma$.
In Fig. 2(e), we plot steady-state population versus the trapping frequency for various LD parameters $\eta$. The population converges for $\eta<0.1$, with the last two curves $\eta=0.1$ and $\eta=0.01$ nearly identical. These curves agree well with Eq. 37 for $\alpha=1$. From this, we can see that getting 99% ground state ( or $\langle n\rangle=10^{-2}$) requires $\nu/\Gamma>5$ for $\eta<0.1$. Typical trapping frequencies in optical tweezers are $\sim 100$ kHz. Because alkali and alkaline-earth linewidths are 5 MHz for the nS to nP transition, sideband cooling cannot prepare ground state atoms.
III.2 Dark sideband cooling
One lesson from the sideband cooling formalism is that the sideband is driven by the gradient of the field, not the field itself. This can be seen by expanding the atom-field interaction as a series to first order in terms of the LD parameter and field gradient
@@ III.2 Dark sideband cooling
This gradient can be a complex phase gradient, for example due to a traveling wave with $e^{ikx}\approx 1+ik\hat{x}$, or an amplitude gradient, for example from a standing wave $\sin(kx)\approx kx$. But that means that a field at the atom equilibrium position is not necessary, and even detrimental as it drives the $n\rightarrow n$ transition with a resulting recoil. The cooling performance of sideband cooling can be improved by adding a second counter-propagating beam with the same polarization to form a standing wave and placing the atom at the node of the field [32]. Although the atom now sees zero intensity at the equilibrium position, the sidebands are still driven. If we place the atom in a standing wave with linear polarization, the field is $\bm{\epsilon}(\hat{x})=\sin(k\hat{x})$. Fig. 2(f) simulates the steady-state population versus atom position with $\nu$ = $-\Delta$. The population reduces by a factor of 5 at the standing wave node. Theoretically, this gets rid of the process in Fig. 2(b) of driving $n$ to $n$, followed by spontaneous emission heating to the sideband. The steady-state solution in Eq. 37 ends up being the same except with $\alpha=0$.
Cooling can be further enhanced by placing the atom in a cavity that can enhance the red sideband [45, 46]. A similar idea could also in principle be obtained by shaping the light mode profile, for example placing the atom at the dark center of a donut beam where a radial gradient still exists, although this has not been theoretically or experimentally investigated.
III.3 Sideband cooling for excited state with different trapping frequency
The Doppler cooling section assumed that the electronic ground and excited states have the same trapping frequency. But typically, the ground and excited states of neutral atoms have different polarizabilities except at “magic wavelengths.” This is due to the excited state being coupled to several higher energy states. The different potentials lead to heating due to the abruptly changing dipole force during absorption or emission, shown in the schematics in Fig. 3. In the extreme case, the excited state is anti-trapped resulting in loss of the atom if the excited lifetime comparable to the oscillation period $1/\nu$.
Because magic wavelengths are not often convenient for optical trapping, neutral atoms more often use two-photon cooling schemes, such as polarization gradient cooling and Raman sideband cooling. However, alkaline-earth atoms contain narrow linewidth transitions that have been used for imaging and cooling [47, 10], even ground state cooling in Ref. [48]. Alkali atoms also contain narrow lines from forbidden transitions like the quadrupole transition of the S to D transition [49] or (n)S to (n+1)P transitions [50].
@@ III.3 Sideband cooling for excited state with different trapping frequency
We now derive the equations for an atom with a ground state frequency $\nu$ and an excited state frequency $\nu_{e}$. The excited state harmonic eigenstates are stretched or compressed relative to the ground state eigenstates. The squeezing operator from quantum optics can therefore be utilized to relate the wavefunctions in the two potentials [52, 53, 54]:
@@
where $|n_{e}\rangle$ is the ground state basis $|m_{e}\rangle$ is the excited state basis. The squeezing operator $\hat{S}(r)$ is
@@
The excited state annihilation operator is related to the ground state annihilation operator by
@@
The harmonic oscillator Hamiltonian is
$H_{\text{ho}}$ $=\nu(\hat{a}^{\dagger}\hat{a})P_{g}+\nu_{e}(\hat{a}_{e}^{\dagger}\hat{a}_{e})P_{e}$
@@
Figure 3: Sideband cooling with different excited $\nu_{e}$ and ground state $\nu$ trapping frequencies. (a) Two regimes, where the excited state is less trapped (left) or more trapped (right) result in extra heating. This heating comes from coupling many different motional states in the quantum picture, or a fluctuating dipole force in the classical picture. For different trapping frequencies, the $n$ to $n-1$ transition can be resonant for only one specific $n$. For $\nu_{e}<\nu$, higher $n$ gradually shift more blue towards heating. For $\nu_{e}>\nu$, higher $n$ gradually shifts more red, leading to cooling of higher order sidebands. (b) Population for a different excited state trapping frequency $\nu_{e}$, plotted for different $\nu/\Gamma$ and $\eta=0.1$ and $\Delta=-\nu$. The magic condition is $\nu_{e}/\nu=1$. (c) Population versus detuning for different excited to ground trapping frequency ratios $\nu_{e}/\nu$. A minimum occurs at $\Delta=-\nu$, but a second minimum occurs at $\Delta=-2\nu$ when $\nu_{e}\neq\nu$. Plot is for $\eta=0.1$, $\nu/\Gamma=2$. The analytical model in Eq. 55 is plotted in a light dashed line. Besides heating, different trapping frequencies also allows the coupling of higher order sidebands, providing new opportunities for cooling.
The transfer rates between the ground states are similar to Eq. 28, where now the squeezing operators are used to modify the excited harmonic states:
@@
Unlike the sideband cooling case with equal trapping frequencies, the sum over the excited states cannot be taken out of the magnitude now because each term does not necessarily have the directional emission terms. There are additional terms now in the first order expansion due to the squeezing operator. To first order in $r$ and $\eta$, the matrix elements are
@@
The first order LD terms still couple $n\rightarrow n\pm 1$. In contrast, the first-order squeezing terms couple $n\rightarrow n\pm 2$, preserving the even and odd symmetry of the wavefunction.
The resonance condition for the light absorption is now determined by the detuning $\Delta+\nu n_{1}-\nu_{e}n_{e}$. We can re-express this in terms of the difference trapping frequency $\Delta\nu=\nu_{e}-\nu$, which gives $\Delta+\nu n_{1}-\nu_{e}n_{e}=(\Delta+n_{e}\Delta\nu)+\nu(n_{1}-n_{e})$. For increasing $n$, the detuning between $n_{1}$ and $n_{e}$ changes by $n_{e}\Delta\nu$. We define the resonance condition for the $n$’th harmonic state as
@@
$\gamma_{n\rightarrow n-1}=\eta^{2}\,n\,A_{\scriptscriptstyle-}.$
@@
$A_{\scriptscriptstyle-}=R({\Delta}_{n}+\nu)+\alpha R({\Delta}_{n}).$
In this context, the rates add incoherently due to the randomized phase acquired during emission in multiple directions.
As illustrated in Fig. 3(a), the $n$-dependent detuning results in the first-order cooling sideband being resonant only for a specific vibrational quantum number $n$. For a less tightly trapped excited state (left side), the cooling light shifts towards heating sidebands for higher $n$ values. Conversely, for a more tightly trapping excited state (right side), lower $n$ values gradually shift towards cooling of higher-order sidebands. This effect leads to instability when $\nu_{e}/\nu<1$, where higher-energy populations heat out of the trap. When $\nu_{e}/\nu>1$, a type of “cap” emerges, with heating occurring below a certain $n$ value and cooling above it. Consequently, this cap effectively traps population below it.
@@
$\gamma_{n\rightarrow n-2}=r^{2}\,n(n-1)\,\,B_{\scriptscriptstyle-}.$
The new rates $B$ are a sum over different excited states, but the interference now matters because the coupling does not involve spontaneous emission to first order,
@@
$B_{\scriptscriptstyle+}=\left|{T}({\Delta}_{n})-{T}({\Delta}_{n}-2\nu)\right|^{2}.$
@@
The steady-state population is a balance between the sideband and squeezing terms. Interestingly, the squeezing does allow additional cooling by making the $n-2$ sideband larger than the $n+2$ sideband. Cooling due to sidebands is $\eta^{2}(A_{-}-A_{+})$, while squeezing cooling is $2r^{2}(B_{-}-B_{+})$. The squeezing dependence is quadratic with $n$ and is more important at larger $n$. Additionally, there is heating from both that goes as $\eta^{2}A_{+}$ and $2r^{2}B_{+}$, respectively. This new coupling does however bring up an interesting question about whether it can provide new opportunities for cooling.
Fig. 3(b) is a master equation simulation of steady-state population versus the excited to ground state trapping frequency ratio $\nu_{e}/\nu$. The magic condition is $\nu_{e}/\nu=1$. Cooling performance drops on both sides, with much more dramatic for a weaker excited state ($\nu_{e}/\nu<1$). Fig. 3(c) is a detuning scan for $\eta=0.1$, $\nu=2\Gamma$, and various $\nu_{e}/\nu$. There is a resonance at $\Delta=-\nu$ due to sideband cooling, but also a second resonance at $\Delta=-2\nu$ due to the squeezing coupling.
For cooling the atom in a ‘non-magic’ trap, a technique known as “Sisyphus Cooling”, has been recently demonstrated in tweezers [10, 55]. In situations where the polarizability of the excited state is greater (or less) than that of the ground state, the atom experiences a deeper (or shallower) trap upon excitation by the cooling laser. This difference enables the implementation of ‘attractive’ (or ‘repulsive’) Sisyphus cooling. In the repulsive Sisyphus cooling method, there is a limit known as the Sisyphus cap to how much the atom can be cooled. However, in the attractive Sisyphus regime, significantly lower temperatures can be achieved within the trap. Ref. [51] proposes that sweeping the cooling laser’s frequency adiabatically could achieve temperatures well below the Sisyphus cap.